OpenStudy (anonymous):

How do I find the inverse and domain of f? f^-1(x)

OpenStudy (anonymous):

substitute the range with the domain of f to get the f^1

OpenStudy (kinggeorge):

Range of \(f\) is the domain of \(f^{-1}\) and the domain of \(f\) is the range of \(f^{-1}\)

OpenStudy (anonymous):

an example would be most helpful

OpenStudy (anonymous):

so the domain of f^-1 would be (∞,∞) and the range would be the same correct? but would f^-1(x)=? The problem is f(x)= In(x-3)

OpenStudy (kinggeorge):

Since the domain for \(f(x) = \ln(x-3)\) is \((3, \infty)\) That would be the range of your inverse function. The domain of the inverse is \((-\infty, \infty)\) however, so you were correct about that.

OpenStudy (anonymous):

do you know the inverse of \[\ln(x)\]?

OpenStudy (anonymous):

I don't know what the inverse of f^-1(x)=

OpenStudy (anonymous):

the inverse of \[\ln(x)\] is \[e^x\] and vice versa

OpenStudy (anonymous):

so \[e^{\ln(x)}=x\] and likewise \[\ln(e^x)=x\]

OpenStudy (anonymous):

if you want the inverse of \[\ln(x-3)\] you can write \[y=\ln(x-3)\] and solve for x \[e^y=x-3\] \[x=e^y+3\] so your inverse is \[f^{-1}(x)=e^x+3\]

OpenStudy (anonymous):

Thank you!